OpenStudy (anonymous):

how to show the work for 9a=6(a+4)

jimthompson5910 (jim_thompson5910):

9a=6(a+4) 9a=6a+24 9a-6a=24 3a=24 a=24/3 a=8

OpenStudy (anonymous):

distribute 6 thru the a + 4. you get 6a+24=9a subtract 6a and you get 3a=24 divide 24/3 and you get A.

OpenStudy (anonymous):

thanks you guys i really appreciate it